OpenStudy (anonymous):

add and simplify 5z/z^2-16 + z/z-4

OpenStudy (lgbasallote):

\[\frac{5z}{z^2 - 16} + \frac{z}{z-4}\] \[\frac{5z}{(z+4)(z-4)} + \frac{z}{z-4}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):